OpenStudy (anonymous):

Future Value of Wise Buying. If a person saves $63 a month by using coupons and doing comparison shopping, (a) What is the amount for year? (b) What would be the future value of this annual amount over 10 years, assuming an interest rate of 4 percent?

OpenStudy (anonymous):

A) Multiply $63 by 12 since there are 12 months in a year to get $756 B) The future value over the next 10 years can be figured out by multiplying $756 * 10 years to get $7560, now multiply that by an interest rate of 4% which would be 0.04 * $7560 = $302.4 (This is your 4 percent interest rate over the next 10 years) So finally, you just add, $302.4 + $7560 = $7862.4

OpenStudy (anonymous):

That's what I thought and someone told me that for part b it was 756 * 1.04 ^ 10= 1,119.06. I am totally confused

OpenStudy (anonymous):

I was just informed by my instructor that the way to solve this is $756(1+0.04)^10=$1,119.06 Future Value $756*1.48=$1,118.88 then I guess you would $756+$1,118.88= $1,867.88

OpenStudy (anonymous):

Ms. Kelly514 this is what I got for this question: $63 x 12 = $756 $756x0.04= $302.4 $302.4+ $7560= $7862.4 $7862.4 is the amount of interest over 10 years